OpenStudy (anonymous):

y+2/7=3x slope and y-intercept?

OpenStudy (anonymous):

slope is m in y=mx+b form b is the y intercept corrisponding to x=0 in this case , your slope is 3 and your y-int is -2/7

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

i know that y=mx+b, but this equation is written differently.

OpenStudy (anonymous):

You just need to solve for y by subtracting 2/7 from both sides. Then it will look just like y=mx+b. y+2/7 = 3x - 2/7 2/7 --------------- y = 3x - 2/7 See it now?